jollo:

Margy is trying to improve her cardio endurance by performing an exercise in which she alternates walking and running 100.0 m each. If Margy is walking at 1.4 m/s and accelerates at 0.20 m/s2 during one of the running portions, what is her final velocity at the end of the 100.0 m? Round your answer to the nearest tenth.

Narad:

The distance is s=100m The initial velocity is u= 1.4m/s The acceleration is a=0.2m/s^2 The final velocity is v=? Apply the equation of motion \[v^2=u^2+2as\] to determine the final velocity v